# A New Causal Rule Learning Approach to Interpretable Estimation of Heterogeneous Treatment Effect

### v2

@@ Appendix A The PDR framework
Figure 7: Overview of the PDR framework of interpretability in different stages where interpretation matters in the data-science life cycle [31]. How CRL enables PDR interpretability is also shown and highlighted in blue blocks.
The PDR framework [31] provides comprehensive guidance and a rich vocabulary for interpretability within the data science life cycle (DSLC) in relevant research. As shown in the lower part (gray texts) of Figure 7, in a typical DSLC, we begin with a specific domain problem for a specific audience and certain data collected to study the problem. Then, in the modeling part, we explore the data and choose certain methods to fit it. After that, we analyze what the model has learned in the post-hoc stage.
Figure 7 also presents the critical concepts in the PDR framework and their relationships in different stages of DSLC where interpretation matters. The framework consists of three desiderata of interpretations, namely predictive accuracy, descriptive accuracy, and relevancy (italic, bold black texts in the upper part). Predictive accuracy evaluates how well the chosen model fits the data. Descriptive accuracy is the degree to which an interpretation objectively captures the relationships learned by the model. Relevancy measures how much insight an interpretation provides into the research question for a particular audience. Interpretation is also classified into model-based and post-hoc interpretability (framed bold black texts). The first interpretability involves using a simpler model to fit the data during the modeling phase to increase descriptive accuracy, in which predictive accuracy generally decreases or remains unchanged. The second one involves using various methods to extract information from a trained model to enhance descriptive accuracy. When choosing a model, there is often a trade-off between both accuracies (dashed line with a two-way arrow), i.e., whether to use a black box model with higher predictive accuracy or a simpler model with higher descriptive accuracy? The answer usually lies in relevancy determined by the problem’s specific context and its audience.
@@ Appendix B Three assumptions of Neyman–Rubin potential outcome framework
Below, we outline the three basic assumptions of the potential outcome framework: 1. Stable Unit Treatment Value Assumption (SUVTA): The potential outcomes of any instance do not vary with the treatments assigned to other instances and there is only one version of each treatment. 2. Ignorability/Unconfoundedness: Treatment assignment is independent of the potential outcomes conditional on the observed covariates, i.e., $A\perp\{Y(1),Y(-1)\}|X=x$. This assumption implies there are no other unmeasured confounding factors, and treatment assignment can be considered as random given the covariates $X$.
@@
[35] gives an equivalent assumption conditioned on the propensity score, i.e., $A\perp\{Y(1),Y(-1)\}|\pi(a,X)$, which is especially useful in high-dimension settings where it is difficult to match samples with the original covariates. Our method is based on this assumption. 3. Positivity/Overlap: $10<Pr(A=1|X=x)<$ 1, i.e., any instance has a positive probability of receiving either treatment given $X=x$. This assumption ensures the existence of samples for both treatment groups so that it is possible to estimate the treatment effect.
@@
RCT setting is usually considered to satisfy these assumptions automatically due to its deliberate design (especially the randomization on treatment assignment). However, observational data has to be pre-processed to make its treatment allocation pseudo-random to satisfy the (Conditional) Unconfoundedness Assumption and to make its sample sizes more balanced between groups to achieve an accepatable or better overlap.
Appendix C Parameters tuning for CRL and baselines
@@
This section outlines several parameters that require tuning when fitting CRL and other baseline models. All methods, including CRL and the baselines, are tuned using 10-fold cross-validation. The performance criteria are mean squared error (MSE) of estimators and ITE-based prediction accuracy for the simulated data and the real-world ASD data, respectively.
C.1 Parameters tuning for CRL
@@
For both the simulated data and the ASD data, we find that CRL’s performance is insensitive to the number of trees, so we set num.trees to 50 in all datasets to reduce computational costs and minimize the generation of irrelevant rules. For both the simulated data and the ASD data, we find that the performance of CRL is insensitive to the number of trees, thus we set to 50 in all data sets to save computation cost and avoid too many fake rules. mtry is tuned using integers from three to nine and four to sixteen, respectively, for simulated and ASD data where nine and sixteen always correspond to the lowest MSE value. For $\lambda$, the one that gives the minimum mean error is chosen from [0.0004,4]. tuning details of all the above parameters are summarized in Table 1. min.node.size appears to be the most sensitive parameter. Its performance is mainly affected by the number of observations and the number of pre-defined true (positive) subgroups. Details for simulated data are shown in Table 2. For ASD data, the best value is 20.
C.2 Parameters tuning for baselines
We directly use the default parameter settings for OWE and PRIM since they have few parameters that require tuning. Both CRL and CRE follow a rule-based workflow and generate rules using forest-based models. Therefore, we set CRE parameters nearly the same with CRL, namely num.trees to 50, and the mean depth of tree to 4. Other CRE parameters are set to default, such as learning rate to 0.01.
Similar to CRL, The best mtry for both CT and CF are still 9 for simulated data and 16 for the ASD data. num.trees of CF is 50 for all data due to its insensitivity. For min.node.size of CF, we find that 20 performs the best in all simulated data and 40 for the ASD data. For CT applied to the ASD data, the optimal min.node.size is 40 while for the simulated data, the value varies and is shown in Table 3.
@@ C.2 Parameters tuning for baselines
Appendix D Descriptive statistics of ASD variables

### v3

@@ Appendix A The PDR framework
Figure 9: Overview of the PDR framework of interpretability in different stages where interpretation matters in the data-science life cycle [yubinPDR2019]. How CRL enables PDR interpretability is also shown and highlighted in blue blocks.
The PDR framework [yubinPDR2019] provides comprehensive guidance and a rich vocabulary for interpretability within the data science life cycle (DSLC) in relevant research [wu2025newcausalrulelearning]. As shown in the lower part (gray texts) of Figure 9, in a typical DSLC, we begin with a specific domain problem for a specific audience and certain data collected to study the problem. Then, in the modeling part, we explore the data and choose certain methods to fit it. After that, we analyze what the model has learned in the post-hoc stage. Figure 9 also presents the critical concepts in the PDR framework and their relationships in different stages of DSLC where interpretation matters. The framework consists of three desiderata of interpretations, namely predictive accuracy, descriptive accuracy, and relevancy (italic, bold black texts in the upper part). Predictive accuracy evaluates how well the chosen model fits the data. Descriptive accuracy is the degree to which an interpretation objectively captures the relationships learned by the model. Relevancy measures how much insight an interpretation provides into the research question for a particular audience. Interpretation is also classified into model-based and post-hoc interpretability (framed bold black texts). The first interpretability involves using a simpler model to fit the data during the modeling phase to increase descriptive accuracy, in which predictive accuracy generally decreases or remains unchanged. The second one involves using various methods to extract information from a trained model to enhance descriptive accuracy. When choosing a model, there is often a trade-off between both accuracies (dashed line with a two-way arrow), i.e., whether to use a black box model with higher predictive accuracy or a simpler model with higher descriptive accuracy? The answer usually lies in relevancy determined by the problem’s specific context and its audience.
@@ Appendix B Three assumptions of Neyman–Rubin potential outcome framework
Below, we outline the three basic assumptions of the potential outcome framework:
@@
1. Stable Unit Treatment Value Assumption (SUVTA): The potential outcomes of any instance do not vary with the treatments assigned to other instances and there is only one version of each treatment.
2. Strong Ignorability/Unconfoundedness: Treatment assignment is independent of the potential outcomes conditional on the observed covariates, i.e., $A\perp\{Y(1),Y(-1)\}|X=x$. This assumption implies there are no other unmeasured confounding factors, and treatment assignment can be considered as random given the covariates $X$.
@@
[rosenbaum1983central] gives an equivalent assumption conditioned on the propensity score, i.e., $A\perp\{Y(1),Y(-1)\}|\pi(a,X)$, which is especially useful in high-dimension data settings where it is difficult to match samples with the original covariates. Our method is based on this assumption.
3. Positivity/Overlap: $0<Pr(A=1|X=x)<1$, i.e., any instance has a positive probability of receiving either treatment given $X=x$. This assumption ensures the existence of samples for both treatment groups so that it is possible to estimate the treatment effect.
@@
RCT setting is usually considered to satisfy these assumptions automatically due to its deliberate design (especially the randomization on treatment assignment). However, observational data have to be pre-processed to make their treatment allocation pseudo-random to satisfy the (Conditional) Unconfoundedness Assumption and to make their sample sizes more balanced between groups to achieve an acceptable or better overlap.
Appendix C Proof of Theorems and Lemmas
C.1 Proof of Theorem 3.1
The proof for Theorem 3.1 can be easily derived from Lemmas C.1 and C.2 given below.
C.2 Additional Lemmas
To prove Theorem 3.1, we need two lemmas.
Lemma C.1.
For any approximation $\hat{\tau}(\mathbf{X})$ of $f_{0}(\mathbf{X})$, the value difference between the $d^{*}(\mathbf{X})$ (induced by $f_{0}(\mathbf{X})$) and the estimated $\hat{d}_{N}$ (induced by $\hat{\tau}(\mathbf{X})$) can be bounded as
$\|V(d^{*})-V(\hat{d}_{N})\|_{2}\leq(\mathbf{E}\|f_{0}(\mathbf{X})-\hat{\tau}(\mathbf{X})\|_{2}^{2})^{\frac{1}{2}}.$ (11)
Furthermore, if we assume gMC holds, the upper bound can be optimized as
$\|V(d^{*})-V(\hat{d}_{N})\|_{2}\leq C^{\prime}(\mathbf{E}\|f_{0}(\mathbf{X})-\hat{\tau}(\mathbf{X})\|_{2}^{2})^{\frac{1+\alpha}{2+\alpha}}.$ (12)
Lemma C.2.
Assume assumptions (1)-(6) hold. For $t>0$, let the tuning parameter $\lambda$ be
$\lambda=16\sqrt{2}t^{2}\sqrt{\frac{\log^{2}(2M)}{N}}.$ (13)
Then for the constant $C_{1}$ which depends on the constants $a,\rho$ and $\sigma^{2}$ and $t$ sufficiently large, with the probability at least $1-\frac{C_{1}}{t^{2}}$, we have
$\frac{2\|\tilde{\mathbf{X}}_{N}\hat{{\beta}}-f_{0}\|_{2}^{2}}{N}+3\lambda\|\hat{{\beta}}-{\beta}^{*}\|_{1}\leq 6\|\tilde{\mathbf{X}}_{N}{\beta}^{*}-f_{0}\|_{2}^{2}/N+\frac{24\lambda^{2}s_{*}}{\phi_{*}^{2}}.$
The proof of the first lemma can be found in [qi2018d]. We give the proof of the Lemma C.2 below.
C.3 Proof of Lemma C.2
Before proof of Lemma C.2, we have
Lemma C.3 (Nemirovski moment inequality).
For $m\geq 1$ and $p\geq e^{m-1}$, we have the following inequality
$\mathbf{E}\max_{1\leq j\leq M}|\sum_{i=1}^{n}\left(\gamma_{j}\left(Z_{i}\right)-\mathbf{E}\gamma_{j}\left(Z_{i}\right)\right)^{m}|\leq[8\log(2M)]^{\frac{m}{2}}\mathbf{E}\left[\max_{1\leq j\leq M}\sum_{i=1}^{n}\gamma_{j}^{2}\left(Z_{i}\right)\right]^{\frac{m}{2}}.$
Proof of Lemma C.2..
With some algebra derivations, we have
$\|\tilde{\mathbf{X}}\hat{\beta}-f_{0}\|_{2}^{2}/N+\lambda\|\hat{\beta}\|_{1}\leq 2\epsilon^{\top}\tilde{\mathbf{X}}(\hat{\beta}-\beta^{*})/N+\lambda\|\beta^{*}\|_{1}+\frac{\|\tilde{\mathbf{X}}\beta^{*}-f_{0}\|_{2}^{2}}{N}.$ (14)
Then for the first term on the right hand, by the Hölder’s inequality we can have
$2|\epsilon^{\top}\tilde{\mathbf{X}}(\hat{\beta}-\beta^{*})/N|\leq(\max_{1\leq j\leq M}2|\epsilon^{\top}\tilde{X}^{(j)}|/N)\|\hat{\beta}-\beta^{*}\|_{1}.$
Let $\lambda_{0}=16\sqrt{2}t^{2}\sqrt{\frac{\log^{2}(2M)}{N}}$ and define a set
$\Lambda:=\{2\max_{1\leq j\leq M}|\epsilon^{\top}\tilde{X}^{(j)}|/N\leq\lambda_{0}\}.$
Through Lemma C.3 with $m=2$, we have
$P(\Lambda^{c})$ $=P(\{2\max_{1\leq j\leq M}|\epsilon^{\top}\tilde{X}^{(j)}|/N\geq\lambda_{0}\})$
$\leq\frac{\mathbf{E}\max_{1\leq j\leq M}4|\epsilon^{\top}\tilde{X}^{(j)}|^{2}/N^{2}}{\lambda_{0}^{2}}$
$\leq\frac{4[8\log(2M)/N]\mathbf{E}[\max_{1\leq j\leq M}\sum_{i=1}^{n}\epsilon_{i}^{2}\tilde{X}_{ij}^{2}/N]}{16\times 32t^{2}\frac{\log^{2}(2M)}{N}}$
$\leq\frac{a^{2}\sum_{i=1}^{N}\mathbf{E}\epsilon_{i}^{2}/N}{16t^{2}\log(2M)}=\frac{a^{2}(4\|\mathbf{X}\gamma_{0}\|_{2}^{2}/N+4\sigma^{2})}{16t^{2}\log(2M)}$
$\leq\frac{a^{2}(\rho\|\gamma_{0}\|_{2}^{2}+\sigma^{2})}{4t^{2}\log(2M)}$
$\leq\frac{C}{t^{2}}.$
Then with probability at least $1-\frac{C}{t^{2}}$, for large enough $t$, we have
$\|\tilde{\mathbf{X}}\hat{\beta}-f_{0}\|_{2}^{2}/N+\lambda\|\hat{\beta}\|_{1}$ $\leq(\max_{1\leq j\leq M}2|\epsilon^{\top}\tilde{X}^{(j)}|/N)\|\hat{\beta}-\beta^{*}\|_{1}+\lambda\|\beta^{*}\|_{1}+\frac{\|\tilde{\mathbf{X}}\beta^{*}-f_{0}\|_{2}^{2}}{N}$
$\leq\lambda_{0}\|\hat{\beta}-\beta^{*}\|_{1}+\lambda\|\beta^{*}\|_{1}+\frac{\|\tilde{\mathbf{X}}\beta^{*}-f_{0}\|_{2}^{2}}{N}.$
Following the same procedure of Lemma 4 in [qi2018d], we have the following conclusion. That is, on the set $\Lambda$, with $\lambda\geq 2\lambda_{0}$,
$4\|\tilde{\mathbf{X}}\hat{\beta}-f_{0}\|_{2}^{2}/N+3\lambda\|\hat{\beta}_{S_{*}^{c}}\|_{1}\leq 5\lambda\|\hat{\beta}_{S_{*}}-\beta_{S_{*}}^{*}\|_{1}+4\|\tilde{\mathbf{X}}\beta^{*}-f_{0}\|_{2}^{2}/N,$
where $S_{*}=\{j:\beta_{j}^{*}\neq 0\}$. Therefore,
$\quad 4\|\tilde{\mathbf{X}}\hat{\beta}-f_{0}\|_{2}^{2}/N+3\lambda\|\hat{\beta}-\beta^{*}\|_{1}$
$=4\|\tilde{\mathbf{X}}(\hat{\beta}-\beta^{*})\|_{2}^{2}/N+3\lambda\|\hat{\beta}_{S_{*}}-\beta_{S_{*}}^{*}\|_{1}+3\lambda\|\hat{\beta}_{S_{*}^{c}}\|_{1}$
$\leq 12\lambda\|\hat{\beta}_{S_{*}}-\beta_{S_{*}}^{*}\|_{1}\leq 12\sqrt{s_{*}}\lambda\|\tilde{\mathbf{X}}(\hat{\beta}-\beta^{*})_{S_{*}}\|_{2}/(\sqrt{N}\phi_{*})$
$\leq\sqrt{s_{*}}\|\tilde{\mathbf{X}}\hat{\beta}-f_{0}+f_{0}-\tilde{\mathbf{X}}\beta^{*}\|_{2}/(\sqrt{N}\phi_{*})$
$\leq 12\sqrt{s_{*}}\|\tilde{\mathbf{X}}\hat{\beta}-f_{0}\|_{2}/(\sqrt{N}\phi_{*})+12\sqrt{s_{*}}\|\tilde{\mathbf{X}}\beta_{*}-f_{0}\|_{2}/(\sqrt{N}\phi_{*})$
$\leq 6\|\tilde{\mathbf{X}}\beta^{*}-f_{0}\|_{2}^{2}/N+\frac{24\lambda^{2}s_{*}}{\phi_{*}^{2}}+2\|\tilde{\mathbf{X}}\hat{\beta}-f_{0}\|_{2}.$
Therefore,
$\frac{2\|\tilde{\mathbf{X}}\hat{{\beta}}-f_{0}\|_{2}^{2}}{N}+3\lambda\|\hat{{\beta}}-{\beta}^{*}\|_{1}\leq 6\|\tilde{\mathbf{X}}{\beta}^{*}-f_{0}\|_{2}^{2}/N+\frac{24\lambda^{2}s_{*}}{\phi_{*}^{2}}.$
Appendix D Parameters tuning for CRL and baselines
@@
This section outlines several parameters that require tuning when fitting CRL and other baseline models. All methods, including CRL and the baselines, are tuned using 10-fold cross-validation. The performance criteria are the mean squared error (MSE) of estimators and ITE-based prediction accuracy for the simulated data and the real-world ASD data, respectively.
D.1 Parameters tuning for CRL
@@
Figure 1: Mean squared error of treatment effect applying CRL on 3 simulated datasets under different num.trees.
Generally speaking, for both the simulated data and the ASD data in this study, we find that CRL’s performance shows little improvement beyond 50 trees (when other parameters are set optimally). We show Figure 1, the results (mean squared error of CRL estimated effect) of tuning num.trees with 3 simulated datasets for your reference. Therefore, we set num.trees to this fixed value to reduce computational costs and minimize the generation of irrelevant rules. mtry is tuned using integers from three to nine and four to sixteen, respectively, for simulated and ASD data, where nine and sixteen always correspond to the lowest MSE value. For $\lambda$, the one that gives the minimum mean error is chosen from [0.0004,4]. Tuning details of all the above parameters are summarized in Table 1. min.node.size appears to be the most sensitive parameter. Its performance is mainly affected by the number of observations and the number of pre-defined true (positive) subgroups. Details for simulated data are shown in Table 2. For ASD data, the best value is 20.
D.2 Parameters tuning for baselines
We directly use the default parameter settings for OWE and PRIM since they have few parameters that require tuning. Both CRL and CRE follow a rule-based workflow and generate rules using forest-based models. Therefore, we set CRE parameters nearly the same as CRL, namely num.trees to 50, and the mean depth of tree to 4. Other CRE parameters are set to default, such as the learning rate to 0.01.
Similar to CRL, the best mtry for both CT and CF are still 9 for simulated data and 16 for the ASD data. num.trees of CF is 50 for all data due to its insensitivity. For min.node.size of CF, we find that 20 performs the best in all simulated data, and 40 for the ASD data. For CT applied to the ASD data, the optimal min.node.size is 40 while for the simulated data, the value varies and is shown in Table 3.
@@ D.2 Parameters tuning for baselines
Appendix E Robustness analyses of CRL with correlated covariates
E.1 Generation of correlated data.
We generate simulated data identical to the data used in our simulation study one but only introduced controlled correlations between specific variables: $r(x_{1},x_{5})=\gamma\times 0.45$, $r(x_{3},x_{4})=\gamma\times 0.34$, $r(x_{1},x_{5})=\gamma\times 0.1$, $r(x_{1},x_{5})=\gamma\times 0.2$. Here, $\gamma\in\{0,1,2\}$ scales the correlation strength between variables, namely no correlation ($\gamma=0$), mild correlation ($\gamma=1$) and strong correlation ($\gamma=2$) between variables. Since $\gamma=0$ corresponds to the original data used in our simulation, we only generate data for $\gamma=1,2$ for all possible combinations of num.obs, num.grp and k. Therefore, for each $\gamma$, we have 54 data sets, and each data set is resampled 100 times to fit CRL and other baseline methods. These datasets are used both for the overall performance evaluation of CRL and the robustness of p-value as well as its interpretation in the rule decomposition analysis.
Table 4: Details of three types of rules.
Rule Type Rule
Data Setting: $\{N=2500,num.grp=1,k=10\}$
true $x_{1}=1$ and $x_{2}=1$
fake without correlation $x_{1}=1$ and $x_{8}>0$
fake with correlation $x_{1}=1$ and $x_{5}>0$
$r(x_{1},x_{5})=0/0.45/0.9$ $x_{1}=0$ and $x_{5}\leq 0$
Data Setting: $\{N=5000,num.grp=3,k=20\}$
true $x_{1}=1$ and $x_{1}+x_{2}+x_{3}\geq 1$
fake without correlation $x_{1}=0$ and $x_{2}=0$ and $x_{8}>0$
fake with correlation $x_{1}+x_{2}+x_{3}\geq 3$ and $x_{4}>0$
$r(x_{3},x_{4})=0/0.34/0.68$ $x_{1}+x_{2}+x_{3}=0$ and $x_{5}\leq 0$
Data Setting: $\{N=10000,num.grp=5,k=30\}$
true $x_{1}+x_{2}\geq 1$ and $x_{3}>5$
fake without correlation $x_{1}+x_{2}=0$ and $x_{8}>0$
fake with correlation $x_{1}+x_{2}\geq 1$ and $x_{5}>0$
$x_{3}>5$ and $x_{4}\leq 0$
E.2 Performance evaluation of CRL under correlation
The results are shown in Figure 2 and 3 for $\gamma=1$ and $\gamma=2$ respectively, which indicates that variable correlation barely harms the performance of CRL, and that the superiority of CRL over other methods in uncorrelated cases remains. Specifically, CRL keeps the lowest MSE (Figure 2(a) and 3(a) ) and the highest MPO (Figure 2(b) and 3(b)) together with CF in all settings for both $\gamma$. Also, CRL has comparable performance to baseline methods with regard to PO (Figure 2(c) and 3(c)) across most settings. In addition, the proportion of fake CF rules filtered out by CRL (Figure 2(d) and 3(d)) remains high, suggesting CRL’s ability to distinguish true rules from fake ones, even when there are redundant variables highly correlated with critical variables.
Figure 2: Performance comparison of CRL (black asterisk) and baseline methods CT (blue circle), CF (green square), OWE (yellow triangle) applied on the correlated ($\gamma=1$) simulated data. All performance metrics are averaged over 100 repetitions for each data set. (a) Mean squared error of treatment effect. (b) Mean potential outcomes. (c) Mean population overlap. (d) Proportion of fake CF rules filtered out by CRL.
Figure 3: Performance comparison of CRL (black asterisk) and baseline methods CT (blue circle), CF (green square), OWE (yellow triangle) applied on the correlated ($\gamma=2$) simulated data. All performance metrics are averaged over 100 repetitions for each data set. (a) Mean squared error of treatment effect. (b) Mean potential outcomes. (c) Mean population overlap. (d) Proportion of fake CF rules filtered out by CRL.
E.3 Validation of rule decomposition interpretation
Three Types of Rules for Testing. The variables generated in this section can be classified into three categories: set of critical variables which defines the true subgroup i.e., $\{x_{1},x_{2},x_{3}\}$, set of redundant variables which are related to certain critical variables when $\gamma\neq 0$ i.e., $\{x_{4},x_{5}\}$ and the set of the rest variables. When a redundant variable, strongly correlated with a critical one, is removed, the resulting $p$-values might be influenced.
To check if this issue matters, we create various rules by selecting variables from the above sets and check how the p-values of interest change in different combinations of uncorrelated/correlated variables. Specifically, we create three types of rules: true rules, fake rules without variable correlation, and fake rules with variable correlation. Details are shown in Table 4. Note that for simplicity, we only test the results of the following three combinations of num.obs, num.grp, and k: $\{N=2500,num.grp=1,k=10\}$ , $\{N=5000,num.grp=3,k=20\}$, $\{N=10000,num.grp=5,k=30\}$. For each data setting and a given $\gamma$, we only use 10 random datasets to obtain the results. Besides, the rules for the same type differ across data settings as the definition of true subgroups differs (See Figure 3 for the definition of true subgroups). Additionally, due to the complex nature of rule definition and varying magnitude of positive treatment effect for settings with 3 and 5 true subgroups, the rules for testing are represented in a way to cover as many true populations as possible to avoid trivial comparison. For example, we use $x_{1}=1$ and $x_{1}+x_{2}+x_{3}\geq 1$ for $num.grp=3$, instead of testing multiple true combinations of $x_{1},x_{2},x_{3}$.
We then perform rule decomposition analysis for the rules shown above and compare the p-values of revised rules to the complete rule. For each revised rule, we compute $\Delta p=p_{revised}-p_{complete}$, i.e., the difference of p-values between the revised rule and the complete rule. In terms of our decomposition analysis, a positive difference ($\Delta p>0$) means the component removed is critical to the complete rule; otherwise, it might be redundant. Therefore, a rule with a redundant variable (or negative $\Delta p$) will not survive in the rule decomposition analysis step and will be considered as not informative enough for further human validation. Since each rule is analyzed on 10 resampled data, we count the number of positive and negative $\Delta p$ respectively and compare these numbers across $\gamma=0,1,2$.
Results. Figure 4,5 and 6 show the number of positive and negative $\Delta p$ under different $\gamma$ for each specific type of rule and data setting. For example, Figure 4 shows the results for the three types of rules of $\{N=2500,num.grp=1,k=10\}$:
  • For the true rule $x_{1}=1$ and $x_{2}=1$ (upper left), we observe that when $\gamma=0$, removing $x_{1}$ (critical variable, blue pillars) results in a positive $\Delta p$ 7 times and a negative $\Delta p$ 3 times out of 10 randomized trials, demonstrating the importance of $x_{1}$. This importance does not decay as $\gamma$ increases, as we observe 10 times of positive $\Delta p$ when $\gamma=1$ and $\gamma=2$. Similarly, removing $x_{2}$ (critical variable, pink pillars) yields 10, 9, and 9 times of positive $\Delta p$. Therefore, this true rule is more likely to survive in our rule analysis step.
  • For the fake rule without correlation (upper right) $x_{1}=1$ and $x_{8}>0$, we find a clear contrast of removing $x_{1}$ against removing $x_{8}$. While $x_{1}$ remains important to the complete rule (10 positive $\Delta p$), $x_{8}$ exhibits irrelevance to the true subgroup with 10 negative $\Delta p$ across all $\gamma$. Hence, in all cases, this fake rule will be filtered out in our rule analysis step.
  • For the fake rule with correlation (bottom) $x_{1}=1$ and $x_{5}>0$, $x_{1}=0$ and $x_{5}\leq 0$, we find the contrast between critical variable and redundant variable remains, except for a few cases (1 and 3) of positive $\Delta p$ when the correlation is strong ($\gamma=2$), suggesting a small chance of falsely regarding $x_{5}$ as a critical variable and the rule survives in our analysis step.
We find nearly identical results for the other two data settings (Figure 5 and 6), where true rules all survive and other fake rules are all filtered out. Empirically, we can tell that although strong correlations may alter the positive $\Delta p$ signal of critical variables (lower left of Figure 4, lower right of Figure 5), the negative $\Delta p$ signal of redundant variables does not diminish dramatically as correlation increases. This is because when the critical variable is removed, the redundant variable can substitute for the removed variable to a certain degree, leading to the uncertainty of $\Delta p$. However, removing the redundant variable always makes the revised rule closer to the true rule and hence a negative $\Delta p$. Even in rare cases where such rules survive, the strong correlation between variables in these rules can easily be identified by human experts and deemed invalid. For instance, $x_{1}=0$ and $x_{5}\leq 0$ in the lower right of Figure 4). Based on the above findings, we conclude that variable correlations do not substantially compromise the overall validity of our rule decomposition analysis.
Figure 4: Results of rule decomposition analysis under different strengths of variable correlations ($\gamma=0,1,2$) for $\{N=2500,num.grp=1,k=10\}$. The pillars show the count of positive (above 0 vertical line) and negative (below 0 vertical line) $\Delta p$ corresponding to the first component removed (blue) and the second component (pink) removed.
Figure 5: Results of rule decomposition analysis under different strengths of variable correlations ($\gamma=0,1,2$) for $\{N=5000,num.grp=3,k=20\}$. The pillars show the count of positive (above 0 vertical line) and negative (below 0 vertical line) $\Delta p$ corresponding to the first component removed (blue) and the second component (pink) removed.
Figure 6: Results of rule decomposition analysis under different strengths of variable correlations ($\gamma=0,1,2$) for $\{N=10000,num.grp=5,k=30\}$. The pillars show the count of positive (above 0 vertical line) and negative (below 0 vertical line) $\Delta p$ corresponding to the first component removed (blue) and the second component (pink) removed.
Appendix F Descriptive statistics of ASD variables